Volcanic activity has remained confined mainly to the upper north to north-eastern area of the lava dome, overlooking the Whites Tuitts and Mosquito Ghauts. Regular rockfalls and several short lived pyroclastic flows have resulted over the past 24 hours. Most of the flows travelled down the north and north-eastern sectors of the dome. In the process the small ash clouds which were generated were all swept slightly north of west over the St. Georges Hill, Richmond Hill and Garibaldi Hill areas out to the sea. Most travelled at altitudes above 4000 feet at moderate speeds so ash precipitation was mostly at sea.
Cloud cover around the summit area of the volcano has been variable for most of the day. Both the southern and north-eastern sides of the dome were observed. During a five hour occupation of the South Soufriere Hill observation post, no rockfall activity was witnessed down the Galways face of the dome. The summit region cleared occasionally, revealing that there has been little or no change to the entire sector of the dome above the Galways soufriere. The face remained very steep, composed mainly of ash and small aggregate-sized material with several larger blocks distributed about the mid to lower flanks. A small spine which was seen at the top two weeks ago still remains intact. Conversely, numerous rockfalls and small pyroclastic flows were witnessed along the northern and north-eastern sectors, suggesting that growth is still occurring in the summit area and on the upper northern and eastern flanks. The steep and noticeably blocky northern flanks remained unchanged since yesterday. A small pyroclastic flow was seen entering Tuitts ghaut.
The pattern of seismicity remained similar to the last few days, with a predominance of rockfall signals and a relatively small quantity of earthquakes. There was 98 rockfalls, slightly more than yesterday, and 3 hybrid earthquakes. There were also 2 long-period earthquakes.
Further COSPEC runs were made today along the south. The results for todays and yesterdays traverses are not yet complete. The results of Thursdays and Fridays traverses are 770 and 510 tons per day respectively.
Observations today indicate that the dome is still growing and is very active at the summit and in a broad sector which includes the upper northern to south eastern flanks. Further dome growth will continue to steepen and destabilise the existing structure. Pyroclastic flows generated on the northern side of the dome may enter the northern Ghauts on a more regular basis during the next few days. Only essential visits should be made to the central corridor and it should not be used as a route to the airport. Large pyroclastic flows could occur with no warning in the Tar River and White River valleys. Nobody should enter zones A and B, and only essential visits should be made to the evacuated zone. Ash masks should be worn when in the ashy areas.